# Improvement on the Linearity Response of PandaX-4T with new Photomultiplier Tubes Bases

### 31 Dec 2023 v1

@@ Abstract
With the expanding reach of physics, xenon-based detectors such as PandaX-4T in the China Jinping Underground Laboratory aim to cover an energy range from sub-keV to multi-MeV. A linear response of the photomultiplier tubes (PMTs) is required for both scintillation and electroluminescence signals. Through a dedicated bench test, we investigated the cause of the non-linear response in the Hamamatsu R11410-23 PMTs used in PandaX-4T. The saturation and suppression of the PMT waveform observed during the commissioning of PandaX-4T were caused by the high-voltage divider base. The bench test data validated the de-saturation algorithm used in the PandaX-4T data analysis. We also confirmed the improvement in linearity of a new PMT base design, which will be used to upgrade the PMT readout system in PandaX-4T.
@@ 1 Introduction
In recent years, the liquid xenon (LXe) time projection chamber (TPC) has become a leading experimental technique for rare event searches in underground laboratories [1, 2]. Traditional dark matter direct detection experiments, such as PandaX-4T [3], XENONnT [4, c2020xenonenergy], and LZ [5], have transitioned to the multi-ton-scale stage and expanded the physics reach beyond dark matter. LXe detectors also explore additional physics, including double beta decay of Xe [6, 7, 8, 9], double electron capture of Xe [10], and solar B neutrino searches using coherently elastic nuclear scattering (CENS) [11, 12]. The energy range of interest spans from sub-keV to a few MeV, necessitating a stringent energy response in terms of linearity and resolution for dark matter and neutrino physics.
PandaX-4T employs two arrays of Hamamatsu R11410-23 photomultiplier tubes (PMTs) [13] to capture the prompt scintillation signals and delayed electroluminescence signals. The photon response of the PMTs determines the detector response. The initial version of the custom-designed high voltage divider and readout circuit base for the PMTs (referred to as PMT base later) in PandaX-4T focused on dark matter detection in the keV energy range [14]. To address the possible radioactivity of capacitors, only three capacitors were used at the last stages of the dynodes. However, this design choice resulted in PMT waveform distortion and a non-linear energy response for signals at the MeV level. To mitigate this effect, a de-saturation algorithm was developed for double beta decay analysis [6].
@@
Moreover, in nearby signals, the latter is likely to be suppressed by the first large signal. The suppression effect is difficult to correct due to the complicated detector response. We design a modern base with five de-saturation capacitors to achieve better detector linearity at the MeV scale. Such a design can meet the requirements of the PandaX-4T and will be installed in the upcoming detector update.
@@ 2 PandaX-4T detector and linearity requirement of PMTs
The number of photoelectrons (PEs) collected by a PMT ranges from one to for an energy range from sub-keV to multiple MeV. For a 2.6 MeV gamma emitted by the Th decay chain, the expected number of electroluminescence photons is over one million in PandaX-4T based on NEST model [17]. The typical quantum efficiency of the PMT is approximately 30% for the 175 nm wavelength of xenon scintillation light. The S2 signal is approximately one million and 350,000 PEs for the top and bottom arrays, respectively. Therefore, the brightest PMT on the top array may receive PEs or more. The number of PEs cannot be reduced by lowering the PMT gain since a high gain is required to maintain a high detection efficiency for a single PE (SPE) [18]. The typical PMT gain is with a high voltage of 1500 V between the photocathode and the anode.
@@ 2.1 Saturation effect of high energy events
A de-saturation algorithm is developed to mitigate PMT waveform distortion. The summed PE value of a distorted waveform no longer reflects the amount of photons collected. The deviation increases as the intensity of the light increases. To address this, we first define a reference waveform by aggregating non-saturated waveforms of the signal of the same event. The reference waveform is calculated for each signal since it contains the timing information of photons arriving at the PMT arrays. In the PandaX-4T commissioning data, the non-saturation waveform is defined as the ones with fewer than 900 PEs and more than 50 PEs. For a typical 2.6 MeV gamma event, multiple (>10) non-saturated waveforms can always be identified. The summation of these waveforms is intended to minimize fluctuations arising from the rising edges of the individual PMT waveforms. The summation also averages out the minor timing profile differences of PMTs at different positions. The reference waveform is then scaled to align the rising edge with the saturated waveforms, as shown in Fig. 2. The optimal scale factor is achieved by minimizing the chi-square values between the two waveforms. The de-saturated charge is the number of PEs of the scaled reference waveform.
The de-saturation algorithm presumes that the rising edge of a waveform remains unaffected even if the later part of the waveform is distorted. The presumption is validated by the improved detector position and energy reconstruction improvement in the PandaX-4T data [6]. The bench test data further confirms it, as we will describe later.
@@ 2.2 Suppression effect of multi-scattering events
We quantify the suppression effect with double-site events of 2.6 MeV gamma in the PandaX-4T commissioning data. The suppression factor is defined as the charge ratio of observed S2 and expected S2 100s away from S2, where the suppression effect becomes unnoticeable with this large time gap. In Fig. 3, PMTs with 100, 600, and 1000 PEs in the S2 signals are selected, and the suppression factor of the second site as a function of the time interval between the second and the first signals is shown. As one can see, even for a moderately large S2 waveform of 600 PE, the second waveform may be suppressed to approximately 90% when the time interval is small. For a large S2 waveform of 1000 PE, the second waveform can be only half the size without suppression. Considering the typical drift velocity of 1.4 mm/s during the PandaX-4T commissioning runs, the "safe" vertical distance should be more than 14 cm for a preceding 1000-PE waveform.
@@ 3 Improved PMT base
The design of PandaX-4T PMT bases emphasizes low radioactivity requirements and tries to limit the number of capacitors, which is usually more radioactive than other components. Besides the capacitor C1 to ground, the original base used two decoupling capacitors at the last stages, which maintains voltage between two dynodes when the electrical pulse is being read out and improves the linearity of the PMT. In the improved design, we add three more decoupling capacitors from DY8 to DY10, and each capacitor is connected to the ground directly, as shown in Fig. 4(b). As we will show later, the additional capacitors improve the dynamic range of the anode significantly. A new model of Knowles capacitor is found with only half the radioactivity of the current 4T capacitor, based on the assay results from two high purity germanium (HPGe) counting stations [19] in CJPL.
@@ 4.1 Bench test setup
To confirm the saturation and suppression symptoms of the original PMT base, validate the de-saturation algorithm independently, and benchmark the performance of the improved base, we set up a dedicated PMT bench test as shown in the Fig. 6. The system consists of four R11410-23 PMTs equipped with different bases, an arbitrary pulse generator (RIGOL DG4162), two blue light LEDs, and a data acquisition (DAQ) system.
@@ 4.2 Extending of linearity range
The saturation is correlated with the total anode charge and not influenced by the waveform shape, as shown in Fig. 7. The horizontal axis represents the number of electrons collected by the PMT anode in response to an LED light signal. The vertical axis represents the ratio of the measured anode charge to the expected anode charge, which is scaled from monitor PMT anode charges. The ratio represents the linearity of PMT response, and any deviation from the unit means non-linearity. If we define the threshold at 90%, PMTs with the original bases show severe saturation at anode charge of and electrons for high and low PMT gains, respectively. For PMTs with improved bases, the threshold corresponds to and electrons. If we convert the anode charge to the number of PEs, the linear response range of the PMT has been extended to approximately 50000 PE with the improved base.
@@
The de-saturation algorithm demonstrated in Fig. 2 is validated with our bench test. Fig. 8 shows the observed charge and corrected charge with de-saturation versus the true charge, which is again calculated from the monitor PMT. When the number of true charges is more than PEs for PMTs with the original base, the measured charge deviates significantly from expectation. The de-saturation correction algorithm can rectify this over an extensive dynamic range. Nevertheless, uncertainties increase with higher charges due to the severity of saturation, where the saturated waveform loses sensitivity to reflect the true charge accurately. Although de-saturation can alleviate the problem, an improved base is a more effective solution.
The PMT suppression is verified qualitatively with our bench test. In the tests, the light intensity and timing profile of the two LEDs were fixed, and the time interval between the emissions of the two LED lights was adjusted to mimic double-site events in the PandaX-4T detector. The time intervals range from 7 s to 100 s. Fig. 9 (a) and (b) illustrate the suppression effect to the second waveform for PMTs with the original bases. In Fig. 9 (a), a relatively small charge of 500 PE in the first waveform still causes noticeable suppression to the second waveform right after it. When the time interval increases to 47 s, the suppression is less serious than the first case. If we increase the first PMT waveform to have more than 20000 PEs, the suppression effect is much more pronounced, even if the time interval increases, as demonstrated in Fig. 9 (b). One can also notice the distortion for the first and second waveforms in Fig. 9 (b). However, with the use of the improved PMT base, the suppression effect is no longer apparent(Fig. 9 (c)).
@@ 5 Summary
To address these issues further, we have redesigned the voltage divider base to include six de-saturation capacitors while ensuring that the overall radioactivity of the base remains unchanged. PMTs with the new base achieve a linearity range of up to 40000 PEs. The suppression effect is no longer visible from our qualitative tests.

### 08 Apr 2024 v2

# Improvement on the Linearity Response of PandaX-4T with new Photomultiplier Tube Bases

@@ Abstract
With the expanding reach of physics, xenon-based detectors such as PandaX-4T in the China Jinping Underground Laboratory aim to cover an energy range from sub-keV to multi-MeV. A linear response of the photomultiplier tubes (PMTs) is required for both scintillation and electroluminescence signals. Through a dedicated bench test, we investigated the cause of the non-linear response in the Hamamatsu R11410-23 PMTs used in PandaX-4T. The saturation and suppression of the PMT waveform observed during the commissioning of PandaX-4T were caused by the high-voltage divider base. The bench test data validated the de-saturation algorithm used in the PandaX-4T data analysis. We also confirmed the improvement in linearity of a new PMT base design with three more low radioactivity capacitors at later dynodes, which will be used to upgrade the PMT readout system in PandaX-4T.
@@ 1 Introduction
In recent years, the liquid xenon (LXe) time projection chamber (TPC) has become a leading experimental technique for rare event searches in underground laboratories [1, 2]. Traditional dark matter direct detection experiments, such as PandaX-4T [3], XENONnT [4], and LZ [5], have transitioned to the multi-ton-scale stage and expanded the physics reach beyond dark matter. LXe detectors also explore additional physics, including double beta decay of Xe [6, 7, 8, 9], double electron capture of Xe [10], and solar B neutrino searches using coherently elastic nuclear scattering (CENS) [11, 12]. The energy range of interest spans from sub-keV to a few MeV, necessitating a stringent energy response in terms of linearity and resolution for dark matter and neutrino physics.
PandaX-4T employs two arrays of Hamamatsu R11410-23 photomultiplier tubes (PMTs) [13] to capture the prompt scintillation signals and delayed electroluminescence signals. The photon response of the PMTs determines the detector response. The initial version of the custom-designed high voltage divider and readout circuit base for the PMTs (referred to as PMT base later) in PandaX-4T focused on dark matter detection in the keV energy range [14]. To address the possible radioactivity of capacitors, only three capacitors were used at the last stages of the dynodes. However, this design choice resulted in PMT waveform distortion and a non-linear energy response for signals at the MeV level. To mitigate this effect, a de-saturation algorithm was developed for double beta decay analysis [6, 15].
@@
Moreover, for multi-site events that deposit energy within a narrow region, the latter signals are likely suppressed by the first large signal. The suppression effect is difficult to correct due to the complicated detector response. We design a modern base with five de-saturation capacitors to achieve better detector linearity at the MeV scale. Such a design can meet the requirements of the PandaX-4T and will be installed in the upcoming detector update.
@@ 2 PandaX-4T detector and linearity requirement of PMTs
The number of photoelectrons (PEs) collected by a PMT ranges from one to for an energy range from sub-keV to multiple MeV. For a 2.6 MeV gamma emitted by the Th decay chain, the expected number of electroluminescence photons is over one million in PandaX-4T based on NEST model [18]. The typical quantum efficiency of the PMT is approximately 30% for the 175 nm wavelength of xenon scintillation light. The S2 signal produces approximately one million PEs in the top array and 350,000 PEs in the bottom array. Therefore, the brightest PMT on the top array may receive PEs or more. The number of PEs cannot be reduced by lowering the PMT gain since a high gain is required to maintain a high detection efficiency for a single PE (SPE) [19]. The typical PMT gain is with a high voltage of 1500 V between the photocathode and the anode.
@@ 2.1 Saturation effect of high energy events
A de-saturation algorithm is developed to mitigate PMT waveform distortion. The summed PE value of a distorted waveform no longer reflects the amount of photons collected. The deviation increases as the intensity of the light increases. To address this, we first define a reference waveform by aggregating non-saturated waveforms of the signal of the same event. The PMTs from the top and bottom array with charges between 50-900 PE are selected as reference PMTs.The reference waveform is calculated for each signal since it contains the timing information of photons arriving at the PMT arrays. The upper threshold of 900 PE for channel waveform selection is optimized by scanning various charge thresholds to minimize the energy resolution at 2.6 MeV from Th-232 calibration events. For a typical 2.6 MeV gamma event, multiple (>10) non-saturated waveforms can always be identified. The summation of these waveforms is intended to minimize fluctuations arising from the rising edges of the individual PMT waveforms. The summation also averages out the minor timing profile differences of PMTs at different positions. The reference waveform is then scaled to align the rising edge with the saturated waveforms, as shown in Fig. 2. The optimal scale factor is achieved by minimizing the chi-square values between the two waveforms. The de-saturated charge is the number of PEs of the scaled reference waveform.
The de-saturation algorithm presumes that the rising edge of a waveform remains unaffected even if the later part of the waveform is distorted. The presumption is validated by the improved detector position and energy reconstruction improvement in the PandaX-4T data [6]. However, this method has a large uncertainty due to the fluctuation of the rising edge of the saturated waveform. The bench test data further confirms it, as we will describe later.
@@ 2.2 Suppression effect of multi-scattering events
We quantify the suppression effect with double-site events of 2.6 MeV gamma in the PandaX-4T commissioning data. The suppression factor is defined as the charge ratio of observed S2 and expected S2 100s away from S2, where the suppression effect becomes unnoticeable with this large time gap. In Fig. 3, PMTs with 100, 600, and 1000 PEs in the S2 signals are selected, correspondingly the actually S2 signals are 1250, 950 and 500 PE. The suppression factor of the second site as a function of the time interval between the second and the first signals is shown. As one can see, even for a moderately large S2 waveform of 600 PE, the second waveform may be suppressed to approximately 90% when the time interval is small. For a large S2 waveform of 1000 PE, the second waveform can be only half the size without suppression. Considering the typical drift velocity of 1.4 mm/s during the PandaX-4T commissioning runs, the "safe" vertical distance should be more than 14 cm for a preceding 1000-PE waveform.
@@ 3 Improved PMT base
The design of PandaX-4T PMT bases emphasizes low radioactivity requirements and tries to limit the number of capacitors, which is usually more radioactive than other components. Besides the capacitor C1 to ground, the original base used two decoupling capacitors at the last stages, which maintains voltage between two dynodes when the electrical pulse is being read out and improves the linearity of the PMT. In the improved design, we add three more decoupling capacitors from DY8 to DY10, and each capacitor is connected to the ground directly, as shown in Fig. 4(b). As we will show later, the additional capacitors improve the dynamic range of the anode significantly. A new model of Knowles capacitor is found with only half the radioactivity of the current 4T capacitor, based on the assay results from two high purity germanium (HPGe) counting stations in CJPL. The radioactivity of the new base is at the same level as the initial one [20].
@@ 4.1 Bench test setup
To confirm the saturation and suppression symptoms of the original PMT base, validate the de-saturation algorithm independently, and benchmark the performance of the improved base, we set up a dedicated PMT bench test as shown in Fig. 6. The system consists of four R11410-23 PMTs equipped with different bases, an arbitrary pulse generator (RIGOL DG4162), two blue light LEDs, and a data acquisition (DAQ) system.
@@ 4.2 Extending of linearity range
The saturation is correlated with the total anode charge and not influenced by the waveform shape, as shown in Fig. 7. The horizontal axis represents the number of electrons collected by the PMT anode in response to an LED light signal. The vertical axis represents the ratio of the measured anode charge to the expected anode charge, which is scaled from monitor PMT anode charges. The ratio represents the linearity of PMT response, and any deviation from the unit means non-linearity. If we define the threshold at 90%, PMTs with the original bases show severe saturation at anode charge of and electrons for high and low PMT gains, respectively. For PMTs with improved bases, the threshold corresponds to and electrons. If we convert the anode charge to the number of PEs, the linear response range of the PMT has been extended to approximately 40000 PE with the improved base.
@@
The de-saturation algorithm demonstrated in Fig. 2 is validated with our bench test. Fig. 8 shows the observed charge and corrected charge with de-saturation versus the true charge, which is again calculated from the monitor PMT. S2-like waveform with a width of several s is used in the test. When the true charge is a few hundred, the Gauss-like shape of the waveform is not valid, resulting in the unexpected tail at the beginning. When the number of true charges is more than PEs for PMTs with the original base, the measured charge deviates significantly from expectation. The de-saturation correction algorithm can rectify this over an extensive dynamic range. Nevertheless, uncertainties increase with higher charges due to the severity of saturation, where the saturated waveform loses sensitivity to reflect the true charge accurately. Although de-saturation can alleviate the problem, an improved base is a more effective solution.
The PMT suppression is verified qualitatively with our bench test. In the tests, the light intensity and timing profile of the two LEDs were fixed, and the time interval between the emissions of the two LED lights was adjusted to mimic double-site events in the PandaX-4T detector. The time intervals range from 7 s to 100 s. Fig. 9 (a) and (b) illustrate the suppression effect to the second waveform for PMTs with the original bases. In Fig. 9 (a), a relatively small charge of 500 PE in the first waveform still causes noticeable suppression to the second waveform right after it. When the time interval increases to 47 s, the suppression is less serious than the first case. If we increase the first PMT waveform to have more than 20000 PEs, the suppression effect is much more pronounced, even if the time interval increases, as demonstrated in Fig. 9 (b). One can also notice the distortion for the first and second waveforms in Fig. 9 (b). However, with the use of the improved PMT base, the suppression effect is no longer apparent (Fig. 9 (c)). The suppression effect is no longer visible from our qualitative tests with a summed charge below 40000 PE.
@@ 5 Summary
To address these issues further, we have redesigned the voltage divider base to include six de-saturation capacitors while ensuring that the overall radioactivity of the base remains unchanged. PMTs with the new base achieve a linearity range of up to 40000 PEs.